OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

Have you put all the answers into:? \[y = mx + b\]

OpenStudy (anonymous):

No im just totally lost.

OpenStudy (anonymous):

solve each equation for y then you will get \[y = mx + b\] which will help you find the right answer

OpenStudy (anonymous):

The slope is m=-6 and the point (-2,-3) is from the line. Hence the equation is : \[y-(-3)=-6(x-(-2))\] or : \[y+3=-6(x+2)\]